OpenStudy (anonymous):

Sixty is 12% of what number? A. 5 B. 7.2 C. 500 D. 720

OpenStudy (dmezzullo):

5

OpenStudy (e.cociuba):

Convert the 12% to its decimal equivalent of 0.12 and use this equation: 0.12 * n = 60 Divide both sides by 0.12: n = 60 / 0.12 = 500 ANSWER: 500

OpenStudy (e.cociuba):

Explain @dmezzullo

OpenStudy (dmezzullo):

60 divided by 12 is 5

OpenStudy (e.cociuba):

Umm no u have to convert the 12% into a decimal equivalent of 0.12 first.

OpenStudy (dmezzullo):

U get the same answer either way.

OpenStudy (e.cociuba):

But u said 5, it's 500

OpenStudy (e.cociuba):

Here's and easier way to understand: Let x = The number we are looking for. 12% of x = 60. 0.12x = 60 x = 60/0.12 x = 500 @dmezzullo @brooklyn16